OpenStudy (anonymous):

The population N of students in a school can be modeled with the function N = 1250 – 20y, where y is the number of years since 2000. Which statement is true? A. The student population is decreasing by 20 students/year. B. The student population is decreasing by 1250 students/year. C. The student population is increasing by 20 students/year. D. The student population is increasing by 1250 students/year.

OpenStudy (anonymous):

Let's plug in integers for y and see what happens. y = 1 => 1250 - 20 = 1230 y = 2 => 1250 - 40 = 1210 y= 3 => 1250 - 60 = 1490 Do you understand it now?

OpenStudy (anonymous):

I don't think I understand this question myself! This is difficult.:'(